B Richards Company does not separate sales and sales taxes in its register.
The register total for December is $            90,300
The sales tax rate is 5%
INSTRUCTIONS:
Prepare the journal entry to record the sales and the related sales taxes
at December 31, 2019.
3/1/2019 Eclipse Corp. sold a piece of equipment for                25,000
Original cost of equipment                 150,000
Accumulated depreciation as of January 1, 2019                 133,000
Depreciation expense per month                     1,000
INSTRUCTIONS: Prepare the journal entries to record the sale of this equipment
on March 1, 2019
On January 1, 2019, Glaser Corp. issued 5 year 6% bonds payable at face value of              800,000
Interest is payable semi-annually on July 1 and January 1.  
INSTRUCTIONS:   Prepare the journal entries to issue the bonds, pay interest and
accrue the interest at year end.
Answer #1

1.

Sales tax = 5%

Total of sales and sales tax = $90,300

Sales tax = 90,300 x 5/105

= $4,300

Sales revenue = 90,300 - 4,300

= $86,000

Journal

December 31, 2019 Accounts receivables 90,300
Sales 86,000
Sales tax payable 4,300

2.

Original cost of equipment = $150,000

Accumulated depreciation as of January 1, 2019 = $133,000

Depreciation expense per month = $1,000

Depreciation expense for January and February = 1,000 x 2

= $2,000

Accumulated depreciation as of March 1, 2019 = $133,000 + 2,000

= $135,000

Book value of equipment = Original cost of equipment - Accumulated depreciation as of March 1, 2019

= 150,000 - 135,000

= $15,000

Sale price of equipment = $25,000

Gain on disposal of equipment =  Sale price of equipment - Book value of equipment

= 25,000 - 15,000

= $10,000

Journal

March 1, 2019 Cash 25,000
Accumulated depreciation - Equipment 135,000
Equipment 150,000
Gain on disposal of equipment 10,000

3.

January 1, 2019 Cash 800,000
Bonds payable 800,000
July 1, 2019 Interest expense 24,000
Cash 24,000
Dec. 31, 2019 Interest expense 24,000
Interest payable 24,000

Semi annual interest payment = 800,000 x 6% x 6/12

= $24,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.